Find the equation to the straight line passing through the following two points (at1,a/t1) and (at2,a/t2)

Grade:10

2 Answers

Aditya Gupta
4 years ago
once again, we know that the two point form of a st line is given as:
y – y1= [(y2 –  y1)/(x2 – x1)]*(x – x1), this line represents the eqn of a line passing through P(x1, y1) and Q(x2, y2).
here, x1= at1, y1= a/t1 and x2= at2, y2= a/t2
so substitute these values in the std form, we get
y – a/t1= [(a/t2 – a/t1)/(at2 – at1)]*(x – at1)
simplifying,
y*t1*t2 + x= a(t1 + t2)
kindly approve :)
Arun
4 years ago
Dear student
 
Use formula
 
y – y1 =[( y2 – y1) /( x2 – x1)] (x – x1)
 
on putting values
 
x + t1 t2 y = a (t1 + t2)